OpenStudy (anonymous):

A ladder 5m long leans against a house, find the rate at which, 1. the top of the ladder is moving downward if its foot is 3m from the house and moving away at the rate of 0.5m/sec 2.the slope of the ladder decreases

OpenStudy (sriram):

x^2 + y^2=25 differentiate wrt t 2x.dx/dt +2y.dy/dt=0 2*3*0.5+2*4*dy/dt=0 dy/dt=-3/8 means coming down by vel. 3/8

OpenStudy (sriram):

wat is d 2nd question exactly?